Given : Vector A = "No. of letters of your First Name" units @ N 75° E Vector B = "No. of letters of your Last Name" units @ N 25° E Vector C = "No. of letters of your Middle Name" N @ 45° east of south Find the and direction of the resultant vector


1
Expert's answer
2021-07-05T08:42:35-0400
"F=(4\cos{75},4\sin{75})+(8\cos{25},8\sin{25})\\+(7\cos{45},-7\sin{45})=(13.2,2.3)\ N"

"F=13.4\ N\\\theta=9.8\degree\text{ N of E}"